OpenStudy (anonymous):

√20:((7y)^(5)) please help me solve this

OpenStudy (anonymous):

it is a bit confusing the way it is written

OpenStudy (anonymous):

lol sorry let me see if i can rewrite it

OpenStudy (anonymous):

\[\sqrt[20]{7y}^{5}\]

OpenStudy (jwheele1):

one sec

OpenStudy (anonymous):

there is nothing much you can do with this. if it is \[\sqrt[20]{(7y)^5}\] you can cancel a 5 and get \(\sqrt[4]{7y}\)

OpenStudy (jwheele1):

\[\sqrt[20]{7y}^{5}\] is the 5 attached to the y or to 7y? or is it outside of the root?

OpenStudy (jwheele1):

I mean....7 to the 5th power is like 16xxx or something. and the 20th root of that is rediculous. there is not at least 20 Y's so the y would stay in the discriminant.

OpenStudy (jwheele1):

It doesnt seem like we have it right for a homework problem.

OpenStudy (jwheele1):

if the 5 is just on the y than there is nothing you can do